OpenStudy (anonymous):

-9sin^2(x) cos(x) is concave up on what interval? *function is already in second derivative form

OpenStudy (anonymous):

again???

OpenStudy (anonymous):

last time you just helped me with concave down

OpenStudy (anonymous):

and i havent been able to figure concave up since

OpenStudy (anonymous):

hold the phone, it is either concave up or concave down. we found it to be concave up where this is positive right? as i recall it was on the interval where cosine is negative, namely \((\frac{\pi}{2},\frac{3\pi}{2})\)

OpenStudy (anonymous):

wait, i never saw (pi/2, 3pi/2) srry can you help me with 1 to 2 more questions pls

OpenStudy (anonymous):

sure if they are quick because i am about to turn in to a pumpkin

OpenStudy (anonymous):

i need to know the x-coordinate of the point of inflection of this

OpenStudy (anonymous):

-9sin^2(x) cos(x)

OpenStudy (anonymous):

that is where the second derivative changes sign. there are two of them, at \(x=\frac{\pi}{2}\) and also at \(x=\frac{3\pi}{2}\)

OpenStudy (anonymous):

oh ok this is the original equation of the function 3cosx−cos^(3)x for 0<x<2π can you let me know x-intercept and local maximum

OpenStudy (anonymous):

pls

OpenStudy (anonymous):

*original function of the derivative

OpenStudy (anonymous):

we already had the derivative, i believe it was \(-3\sin^3(x)\) right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

this changes sign when \(\sin(x)\) does, at \(x=0\) and \(x=\pi\)

OpenStudy (anonymous):

since \(x=0\) is the endpoint of the interval, you only have one local max at \(x=\pi\)

OpenStudy (anonymous):

scratch that, \(x=\pi\) is a local minimum, not a local max sorry

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

to summerize, your funciton is decreasing on \((0, \pi)\) and increasing on \((\pi,2\pi)\) local min at \(x=\pi\)

OpenStudy (anonymous):

and we already took care of concave up and down etc

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

would that make my local max 2pi?